OpenStudy (anonymous):

Subtract and Simplify : (5x+6)/(x^2-1)-(1)/(x^2-1)

OpenStudy (anonymous):

Denominators are same so you can directly subtract numerator..

OpenStudy (anonymous):

\[\frac{5x + 6}{x^2 - 1} - \frac{1}{x^2 - 1} \implies \frac{5x + 6 - 1}{(x+1)(x-1)} \implies \frac{5 (x+1)}{(x+1)(x-1)}\] After cancelling what are you left with ??

OpenStudy (anonymous):

@jchavez96 can you cancel any factor in numerator and denominator that you are seeing as same??